A financial analyst has found out that p olicyholders are four times as likely to file two
claims as to file four claims. If the numb er of claims filed has a p osson distribution,
what is the variance of the numb er of claims filed.
1
Expert's answer
2020-05-13T19:58:06-0400

Let X be the random number of claims filed by policyholders. We know that the variance of a Poisson distribution is λ. But we have 4P(X = 4) = P(X = 2) that is 4e (λ4/4!)=e2 /2!) ⇒4λ2/4!=1/2 ⇒ λ2 = 3⇐⇒ λ =1.732 hence the variance of X is equal to 1.732